OpenStudy (anonymous):

Addition Equations? Problem is 5+r+1.435=8.435

OpenStudy (anonymous):

so solve for r?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

and then see if it's correct

OpenStudy (anonymous):

8.435 - 1.435 - 4 = r

OpenStudy (anonymous):

- 5*

OpenStudy (anonymous):

r = 2

OpenStudy (anonymous):

thanks very much!! Math is preeeety tuff

OpenStudy (anonymous):

not really. Dont worry you have a long way ahead of you. We're all still learning

OpenStudy (anonymous):

yea exactly

OpenStudy (anonymous):

well if you have any other questions, i'm right here :)

OpenStudy (anonymous):

ok I'll have a few more in a bit

OpenStudy (anonymous):

alright

OpenStudy (anonymous):

ok stll addition eqautions......................... 9.25+x+1.5=12

OpenStudy (anonymous):

alright same thing make the variable by itself

OpenStudy (anonymous):

try it

OpenStudy (anonymous):

so 9.25-1.5?

OpenStudy (anonymous):

nope

OpenStudy (anonymous):

oh i think it could be 12-1.5?

OpenStudy (anonymous):

when you move numbers to the other side you subtract them on the side they were on and that is part of it now get the other number to that side

OpenStudy (anonymous):

1.3? is that the answer?

OpenStudy (anonymous):

its 1.25

OpenStudy (anonymous):

12 - 9.25 - 1.5 = 1/25

OpenStudy (anonymous):

1.25*

OpenStudy (anonymous):

ok well i think i got it from here

OpenStudy (anonymous):

alrii